# OTESGN:Optimal Transport Enhanced Syntactic-Semantic Graph Networks for Aspect-Based Sentiment Analysis

### v1

@@ Abstract
Aspect-based sentiment analysis (ABSA) aims to identify aspect terms and determine their sentiment polarity. While dependency trees combined with contextual semantics effectively identify aspect sentiment, existing methods relying on syntax trees and aspect-aware attention struggle to model complex semantic relationships. Their dependence on linear dot-product features fails to capture nonlinear associations, allowing noisy similarity from irrelevant words to obscure key opinion terms. Motivated by Differentiable Optimal Matching, we propose the Optimal Transport Enhanced Syntactic-Semantic Graph Network (OTESGN), which introduces a Syntactic-Semantic Collaborative Attention. It comprises a Syntactic Graph-Aware Attention for mining latent syntactic dependencies and modeling global syntactic topology, as well as a Semantic Optimal Transport Attention designed to uncover fine-grained semantic alignments amidst textual noise, thereby accurately capturing sentiment signals obscured by irrelevant tokens. A Adaptive Attention Fusion module integrates these heterogeneous features, and contrastive regularization further improves robustness. Experiments demonstrate that OTESGN achieves state-of-the-art results, outperforming previous best models by +1.01% F1 on Twitter and +1.30% F1 on Laptop14 benchmarks. Ablative studies and visual analyses corroborate its efficacy in precise localization of opinion words and noise resistance.
@@
Optimal Transport, Aspect-based Sentiment Analysis, Graph Neural Networks, Dependency Tree
@@ I Introduction
Aspect-based Sentiment Analysis (ABSA) aims to identify specific aspect terms within a text and determine their corresponding sentiment polarity (e.g., positive, negative, or neutral). A key challenge lies in effectively modeling the semantic dependencies between aspect terms and their relevant contextual cues. For example, as illustrated in Fig 1, ABSA needs to recognize the aspect terms "performance" and "cooling", and precisely associate "performance" with "quite good" and "cooling" with "not keep up"—determining a positive sentiment for "performance" and an indirectly expressed negative sentiment for "cooling".
Figure 1: Dependency tree example constructed with Stanford CoreNLP, featuring dual aspects ("performance" and "cooling") exhibiting divergent sentiment polarities.
@@
To enhance the syntactic modeling capability of attention mechanisms, researchers have proposed various solutions. For instance, the authors of [5] introduced a multi-channel syntax-enhanced attention mechanism by incorporating syntactic relative distance (SRD) derived from the shortest dependency path to suppress syntactically irrelevant noise. On the other hand, the paper [6] modeled syntactic and semantic dependencies between words using dependency graphs. The introduction of Graph Neural Networks (GNNs) represents another major line of research. The pioneering work in [7] applied Graph Convolutional Networks (GCNs) and Graph Attention Networks (GATs) [8] to syntactic information integration, stimulating numerous subsequent studies based on GCNs [9] and GATs [10]. Notably, the paper [11] proposed a dual-channel Transformer-enhanced architecture with gated attention, while the authors of [12] constructed aspect-opinion dependency graphs using biaffine attention. The underlying principle unifying these approaches lies in applying differentiated weighting to dependency neighbor nodes within the graph, effectively suppressing noise and improving representation quality.
@@
To tackle the outlined challenges, we introduce the Optimal Transport Enhanced Syntactic-Semantic Graph Network (OTESGN), a novel approach that implicitly models word associations via semantic distribution matching. OTESGN leverages BERT to produce context-aware word embeddings and employs dependency syntax to construct a mask matrix for enhanced structural support. A Syntactic-Semantic Collaborative Attention (SSCA) integrates local and global semantics through two components: Syntactic Graph-Aware Attention (SGAA), which captures global dependencies using syntax-guided self-attention, and Semantic Optimal Transport Attention (SOTA), which computes directed aspect-opinion word associations by solving an optimal transport coupling matrix with Sinkhorn iteration, using the inverse of cosine similarity as the transport cost. A Adaptive Attention Fusion (AAF) module dynamically balances these channels, while multi-layer residual connections update neighborhood features iteratively, with contrastive learning enhancing robustness.Experiments across three benchmark datasets show OTESGN significantly outperforms baselines in accuracy and F1 score, validating its effectiveness in aligning aspect and opinion words and its robustness to noise.
@@
  • We introduce a Syntactic-Semantic Collaborative Attention (SSCA) that combines Syntactic Graph-Aware Attention (SGAA) and Semantic Optimal Transport Attention (SOTA), with a Adaptive Attention Fusion (AAF) module to dynamically integrate features and contrastive regularization to enhance robustness.
@@ II Related work
Figure 2: Overall Architecture of the OTESGN Framework. Bi-Orthogonal Attention comprises two principal channels: SGAA for structural perception and SOTA for semantic transmission, with adaptive fusion mediated by the AAF.
@@ III-B1 Syntactic Graph-Aware Attention
Figure 3: Semantic Distance Modeling for Optimal Transport.
@@ III-B2 Semantic Optimal Transport Attention
Figure 4: Solving Semantic Optimal Transport in the embedding space, where the aspect term is optimally transported toward high-weight context words to achieve fine-grained semantic alignment.
@@ IV-A Datasets
The sample distribution and DD statistics across datasets are illustrated in Fig 5 and Table 1. As shown in Table 1, the average DD values for Laptop14 (3.21) and Rest14 (3.17) are relatively close, while Twitter exhibits a slightly higher average of 3.41. This indicates notable linguistic variation across domains: the informal expressions and complex metaphors common in Twitter data lead to increased syntactic complexity, whereas the more formal and domain-specific terminology in Laptop14 and Rest14 contributes to more uniform syntactic structures. These findings suggest that robust models must be resilient to noisy and diverse linguistic patterns.
@@ IV-D Main Results
TABLE II: Comparison of OTESGN with Baseline Models on Rest14, Laptop14, and Twitter Datasets. Best results are colored in first, second, third (Unit:%).
@@
TABLE III: Ablation Study of OTESGN on Rest14, Laptop14, and Twitter Datasets. “w/o SM”, “w/o GA”, “w/o OT”, “w/o GA+OT”, and “w/o CL” denote removing the Syntactic Mask, SGAA, SOTA, both SGAA & SOTA, and Contrastive Learning, respectively (Unit:%).
@@
Figure 6: The evolution of the learnable parameter $\beta$ across training epochs begins to converge to a stable value from the 4th epoch.
@@ IV-F Case Study
Comprehensive case analysis of misclassified samples (Table IV) reveals the primary error pattern based on the confusion matrix Fig 7, the primary error pattern of the OTESGN model lies in misclassifying positive and negative sentiments as neutral. This is mainly attributed to the absence of explicit sentiment markers and the implicit nature of deep semantic associations in the input. As illustrated in Cases 1 and 2, the lack of directly expressed polarity words leads the model to adopt a more conservative, neutral prediction. Additionally, we observe instances of negative sentiments being misclassified as positive, such as in Case 3. This indicates that when subjective opinions are expressed through discontinuous or nested structures, the model struggles to capture the underlying sentiment reversal effectively. These findings highlight current limitations in implicit sentiment recognition and complex syntactic parsing. Future work could explore the integration of context-aware mechanisms or enhanced syntactic dependency modeling to address these challenges.
@@ IV-G Visualization Analysis
As illustrated in the attention heatmap in Fig 8, the structure-aware graph attention effectively focuses on words syntactically close and semantically related to the aspect term "Sarah Palin" such as "disinformation" (weight=0.27) and "belittling" (weight=0.19), demonstrating strong structural guidance.
Figure 8: Word-Level Attention Heatmap.
Figure 9: Effect of Hyperparameter Eps on OT Attention Scores.
@@
Figure 10: Dropout Regularization and OT Attention Dynamics Across Network Layers.
@@ IV-H Hyperparameter Sensitivity Analysis
This paper proposes a novel OTESGN model. By systematically analyzing domain distribution differences in benchmark datasets and building upon prior research, it innovatively integrates optimal transport theory with syntactic-semantic graphs, achieving optimal semantic association between aspect and opinion terms while effectively reducing noise interference. First, a syntactic mask matrix is constructed based on dependency parsing relationships. Then, a SSCA comprising SGAA and SOTA is designed for global semantic modeling and local optimal semantic alignment. Subsequently, a AAF module dynamically adjusts weights, and contrastive learning is introduced to enhance model robustness. Experiments demonstrate OTESGN’s significant performance advantages across three benchmark datasets, with extensive experimental results validating the effectiveness of the proposed approach.
@@ VI Limitations
On Rest14, OTESGN is competitive but does not surpass the strongest baseline. Likely factors include a stronger reliance on explicit syntax under standardized reviews and occasional structural mismatch from the dual-channel attention. We also observe failures on instances with implicit polarity. Another limitation is the moderate overhead of the OT solver relative to dot-product attention; we mitigate this by tuning the entropy regularization coefficient $\epsilon^{k}$ to balance coupling sharpness and convergence. Future work will explore adaptive syntax extraction and the integration of event/knowledge priors to improve robustness.

### v2

# OTESGN: Optimal Transport-Enhanced Syntactic-Semantic Graph Networks for Aspect-Based Sentiment Analysis

@@ Abstract
Aspect-based sentiment analysis (ABSA) aims to identify aspect terms and determine their sentiment polarity. While dependency trees combined with contextual semantics provide structural cues, existing approaches often rely on dot-product similarity and fixed graphs, which limit their ability to capture nonlinear associations and adapt to noisy contexts. To address these limitations, we propose the Optimal Transport-Enhanced Syntactic-Semantic Graph Network (OTESGN), a model that jointly integrates structural and distributional signals. Specifically, a Syntactic Graph-Aware Attention module models global dependencies with syntax-guided masking, while a Semantic Optimal Transport Attention module formulates aspect-opinion association as a distribution matching problem solved via the Sinkhorn algorithm. An Adaptive Attention Fusion mechanism balances heterogeneous features, and contrastive regularization enhances robustness. Extensive experiments on three benchmark datasets (Rest14, Laptop14, and Twitter) demonstrate that OTESGN delivers state-of-the-art performance. Notably, it surpasses competitive baselines by up to +1.30 Macro-F1 on Laptop14 and +1.01 on Twitter. Ablation studies and visualization analyses further highlight OTESGN’s ability to capture fine-grained sentiment associations and suppress noise from irrelevant context.
@@
Aspect-based Sentiment Analysis, Optimal Transport, Graph Neural Networks, Dependency Tree
@@ I Introduction
Aspect-based Sentiment Analysis (ABSA) aims to identify specific aspect terms within a text and determine their corresponding sentiment polarity (e.g., positive, negative, or neutral). A key challenge lies in effectively modeling the semantic dependencies between aspect terms and their relevant contextual cues. For example, as illustrated in Fig. 1, ABSA needs to recognize the aspect terms "performance" and "cooling", and precisely associate "performance" with "quite good" and "cooling" with "not keep up"—determining a positive sentiment for "performance" and an indirectly expressed negative sentiment for "cooling".
Figure 1: Dependency tree example constructed with Stanford CoreNLP, featuring dual aspects ("performance" and "cooling") exhibiting divergent sentiment polarities
@@
To enhance the syntactic modeling capability of attention mechanisms, researchers have proposed various solutions. For instance, the authors of [5] introduced a multi-channel syntax-enhanced attention mechanism by incorporating syntactic relative distance (SRD) derived from the shortest dependency path to suppress syntactically irrelevant noise. On the other hand, the paper [6] modeled syntactic and semantic dependencies between words using dependency graphs. The introduction of Graph Neural Networks (GNNs) represents another major line of research. The pioneering work in [7] applied Graph Convolutional Networks (GCNs) and Graph Attention Networks (GATs) [8] to syntactic information integration, stimulating numerous subsequent studies based on GCNs [9] and GATs [10]. Notably, the paper [11] proposes a dual-channel Transformer-enhanced architecture with gated attention, while the authors of [12] constructed aspect-opinion dependency graphs using biaffine attention. The underlying principle unifying these approaches lies in applying differentiated weighting to dependency neighbor nodes within the graph, effectively suppressing noise and improving representation quality.
@@
To tackle the outlined challenges, we introduce the Optimal Transport-Enhanced Syntactic-Semantic Graph Network (OTESGN), a novel approach that implicitly models word associations via semantic distribution matching. OTESGN leverages BERT to produce context-aware word embeddings and employs dependency syntax to construct a mask matrix for enhanced structural support. A Syntactic-Semantic Collaborative Attention (SSCA) integrates local and global semantics through two components: Syntactic Graph-Aware Attention (SGAA), which captures global dependencies using syntax-guided self-attention, and Semantic Optimal Transport Attention (SOTA), which computes directed aspect-opinion word associations by solving an optimal transport coupling matrix with Sinkhorn iteration, using the inverse of cosine similarity as the transport cost. An Adaptive Attention Fusion (AAF) module dynamically balances these channels, while multi-layer residual connections update neighborhood features iteratively, with contrastive learning enhancing robustness.Experiments across three benchmark datasets show OTESGN significantly outperforms baselines in accuracy and F1 score, validating its effectiveness in aligning aspect and opinion words and its robustness to noise.
@@
  • We design a Syntactic-Semantic Collaborative Attention (SSCA) that combines Syntactic Graph-Aware Attention (SGAA) and Semantic Optimal Transport Attention (SOTA), with an Adaptive Attention Fusion (AAF) module to dynamically integrate features and Contrastive Regularization to enhance robustness.
@@ II Related work
Figure 2: Overall Architecture of the OTESGN Framework. Bi-Orthogonal Attention comprises two principal channels: SGAA for structural perception and SOTA for semantic transmission, with adaptive fusion mediated by the AAF
@@ III-B1 Syntactic Graph-Aware Attention
Figure 3: Semantic Distance Modeling for Optimal Transport
@@ III-B2 Semantic Optimal Transport Attention
Figure 4: Solving Semantic Optimal Transport in the embedding space, where the aspect term is optimally transported toward high-weight context words to achieve fine-grained semantic alignment
@@ IV-A Datasets
The sample distribution and DD statistics across datasets are illustrated in Fig. 5 and Table 1. As shown in Table 1, the average DD values for Laptop14 (3.21) and Rest14 (3.17) are relatively close, while Twitter exhibits a slightly higher average of 3.41. This indicates notable linguistic variation across domains: the informal expressions and complex metaphors common in Twitter data lead to increased syntactic complexity, whereas the more formal and domain-specific terminology in Laptop14 and Rest14 contributes to more uniform syntactic structures. These findings suggest that robust models must be resilient to noisy and diverse linguistic patterns.
@@ IV-D Main Results
TABLE II: Comparison of OTESGN with Baseline Models on Rest14, Laptop14, and Twitter Datasets. Best results are colored in first, second, third (Unit:%)
@@
TABLE III: Ablation Study of OTESGN on Rest14, Laptop14, and Twitter Datasets. “w/o SM”, “w/o GA”, “w/o OT”, “w/o GA+OT”, and “w/o CL” denote removing the Syntactic Mask, SGAA, SOTA, both SGAA & SOTA, and Contrastive Learning, respectively (Unit:%)
@@
Figure 6: The evolution of the learnable parameter $\beta$ across training epochs begins to converge to a stable value from the 4th epoch
@@ IV-F Case Study
Comprehensive case analysis of misclassified samples (Table IV) reveals the primary error pattern based on the confusion matrix Fig. 7, the primary error pattern of the OTESGN model lies in misclassifying positive and negative sentiments as neutral. This is mainly attributed to the absence of explicit sentiment markers and the implicit nature of deep semantic associations in the input. As illustrated in Cases 1 and 2, the lack of directly expressed polarity words leads the model to adopt a more conservative, neutral prediction. Additionally, we observe instances of negative sentiments being misclassified as positive, such as in Case 3. This indicates that when subjective opinions are expressed through discontinuous or nested structures, the model struggles to capture the underlying sentiment reversal effectively. These findings highlight current limitations in implicit sentiment recognition and complex syntactic parsing. Future work could explore the integration of context-aware mechanisms or enhanced syntactic dependency modeling to address these challenges.
@@ IV-G Visualization Analysis
As illustrated in the attention heatmap in Fig. 8, the structure-aware graph attention effectively focuses on words syntactically close and semantically related to the aspect term "Sarah Palin" such as "disinformation" (weight=0.27) and "belittling" (weight=0.19), demonstrating strong structural guidance.
Figure 8: Word-Level Attention Heatmap
Figure 9: Effect of Hyperparameter Eps on OT Attention Scores
@@
Figure 10: Dropout Regularization and OT Attention Dynamics Across Network Layers
@@ V Conclusion
This paper presented OTESGN, an Optimal Transport-Enhanced Syntactic–Semantic Graph Network for aspect-based sentiment analysis. By integrating syntactic constraints with distributional alignment, the model captures both global dependencies and fine-grained semantic associations. Specifically, Syntactic Graph-Aware Attention encodes structural relations, while Semantic Optimal Transport Attention formulates aspect–opinion matching as a distribution problem solved via the Sinkhorn algorithm. An Adaptive Attention Fusion mechanism balances these signals, and contrastive regularization improves robustness. Experiments on Rest14, Laptop14, and Twitter show that OTESGN consistently outperforms or matches state-of-the-art baselines, with notable gains on Laptop14 and Twitter. Ablation and visualization further highlight the role of optimal transport in enhancing alignment and mitigating noise.
@@ VI Limitations
On Rest14, OTESGN is competitive but does not surpass the strongest baseline. Likely factors include a stronger reliance on explicit syntax under standardized reviews and occasional structural mismatches from the dual-channel attention. Additionally, we observe errors in cases involving implicit polarity, where subtle sentiment cues are not explicitly expressed in the text. Another limitation is the moderate overhead of the OT solver relative to dot-product attention; we mitigate this by tuning the entropy regularization coefficient $\epsilon^{k}$ to balance coupling sharpness and convergence. Future work will explore adaptive syntax extraction and the integration of event/knowledge priors to further improve robustness.